# AI on My Shoulder: Supporting Emotional Labor in Front-Office Roles with an LLM-based Empathetic Coworker

### v1

@@ 6.1. Utility: Functions of Embedding Empathetic AI into Uncivil Conversations
We found that Emo-Label mirrors the function of some coworkers who simply acknowledge the emotional downsides of the job, without actually providing a way forward. P20 called out the desensitization of his coworkers, “I wouldn’t say they’d be helpful, they see it like it’s a part of my job.” By contrast, participants could identify that Emo-Reframe uniquely varied from simple sentiment classification.
@@ 6.1.1. Empathetic messages help CSRs avoid negative thinking traps
In the face of client incivility, a CSR is likely to feel negatively about the situation. Participants reported learning to dissociate from emotions while on task, but the overwhelming volume of client engagement can make complete dissociation challenging (P01, P04, P05, P06, P08, P14, P18, P19, P20). However, as part of their role, a CSR needs to suppress their emotional impulse. P13 described her internal method, “whatever your reaction is, you have to keep that in your head and then decide how you’re going to respond professionally.” While Surface Acting has its benefits (Grandey, 2000), only changing responses at a superficial level contributes to dissatisfaction at work (Côté and Morgan, 2002). Instead, in line with recommendations from emotion literature (Grandey, 2000), we found that Emo-Reframe enabled Deep Acting — adjusting how participants cognitively evaluated their experience.
@@ 6.1.2. Empathetic messages give CSRs an opportunity to recenter their problem solving
Other participants echoed P07’s experience. Essentially, Emo-Reframe helped the participants center their thoughts and continue working towards resolutions. P01 particularly drew a parallel between Emo-Reframe and a coach who would help her maintain “momentum.” Much like how Info-Guide had an affective purpose, participants perceived Emo-Reframe to have a procedural purpose that helped them complete their task. Pro-Pilot’s message drew their attention to their objectives and motivated them to work towards it. Note, we designed the empathetic messages with examples of empathy that were high in specificity (Sharma et al., 2023). Participants’ descriptions indicated regaining confidence and psychological safety due to Emo-Reframe. P05 stated, “it feels validating that you’re not necessarily making a mistake.” As a result, participants envisioned their job proficiency improving with empathetic support embedded in the messages.
@@ 6.1.3. Empathetic messages humanize the client
By reading the Emo-Reframe panel, participants were no longer speculating the perspective of the client. P03 actually became more observant of the client and started modifying his language to be more accommodating. Participants were more willing to service the clients as “someone alive” (P17). The assistance of Emo-Reframe helped the participants rejuvenate their interest in the client’s emotional state. Not only was this meaningful to their personal wellbeing, but also enables them to build longer-term relationships with clients.
@@ 6.2. Complementarity: Situations where Empathetic AI can Stand-In for Coworkers
Organizations employ multiple CSRs to handle a large volume of client interactions. Regardless of the peculiarity in each client, CSRs together form a “united front” (P04). Coworkers play a key role in supporting CSRs and their empathy is fundamental to mitigating the pressures of client incivility (Ho and Gupta, 2014; Sakurai and Jex, 2012). After interacting with Pro-Pilot, our participants compared and contrasted its function with that of their coworkers’. Based on these reflections, we identified two key gaps in existing coworking paradigms of front-office work that emotional support from Pro-Pilot can address.
@@ 6.2.1. On-task emotional AI can reduce burden in coworker support
Interactions with clients are nuanced. When the conversation reaches a deadlock, a CSR might want to reach out to their coworkers. However, in cases of incivility, our participants brought to light a double-sided burden problem. Incivility leads to exhaustion (Brotheridge and Grandey, 2002) and can hinder a CSRs ability to seek support. Consider P01’s example, “It’s just hard to bring them into my conversations when I’ll have to explain to them like what is going on.” Support requires some disclosure, but disclosing the complexities of a situation can be burdensome (Bazarova, 2015). Meanwhile, it can also be burdensome for the coworker to be emotionally available for the problem (Fadden et al., 1987). Since the CSR is likely to reach out to a coworker who shares the client, that coworker might be more focused on the task than emotions. By contrast, they viewed Pro-Pilot as a tool that can be present throughout their task, observe emotional duress, and provide support immediately. P06 captures the unique opportunity of Pro-Pilot’s emotional support:
The pro-pilot was kind of there with me in the moment, whereas I can’t really have those conversations while I’m trying to have a conversation with the customer. I’ve never had a little emotional support buddy, like that before.” — P06
A CSR might put a client on hold to consult a coworker about procedure, but emotional aspects were discussed after the interaction. Instead, “pro-pilot is an immediate solution” (P10) for managing emotions of the task. Our lexical evaluation (Section 4.2.2) also confirmed that Pro-Pilot was capable of adapting to the specifics of the situation better than human coworkers. Another value proposition of Pro-Pilot was that it reduces the need for emotional oversight. P08, who now supervises other CSRs as well, imagined that Pro-Pilot would take away the need to “always be on edge, or a constant vigilance, to make sure that my agents … were (not) being insensitive to the client’s question.” With Pro-Pilot CSRs stand to reprioritize which emotional discussions they should have with their coworkers.
@@ 6.2.2. Simulating incivility with emotional support can train CSR to use ER
Beyond on-boarding manuals, many CSRs learn on-the-job through other more experienced coworkers (Shteynberg and Apfelbaum, 2013). This learning is hampered because of conflicts that arise among coworkers. If one CSR fails at a task, it is typically escalated to their coworkers. Our participants pointed out that the increased workload of the coworker makes them less inclined to teach the CSR better approaches. This arrangement can strain the social ties between front-office workers. “At work, there are no real friends,” said P04. P01 found herself unable to consult coworkers for her emotional needs because they were not close. P07 even experienced being put down by her coworkers in front of clients. These tensions reduce the ability of early career CSRs to learn from exemplar coworkers. Our study recruited CSRs with at least 2 years of experience. Several expressed their desire to have had a tool like Pro-Pilot for training them early.
When they face stuff like this they tend to break down. They need people around them to tell them stuff like what pro-pilot was doing for me.” — P20
In line with P20’s quote above, P13 — another participant whose role includes supervision — stated that exercising client interactions with Emo-Reframe could help CSRs to develop the ”mindset” needed for client interactions. “ Some of those live reminders, could be really beneficial for folks, who are less practiced or just have more trouble with that kind of emotional regulation,” said P13. Training up the CSRs was seen as an important way to make participants self-reliant. Recent literature shows that workers can outperform their personality based stereotype if they follow healthier behavior patterns (Das Swain et al., 2019b). P18 wants to handle client incivility on her own. With Pro-Pilot she felt that her “escalation techniques” would no longer be scrutinized. Some participants saw themselves adapting Pro-Pilot into an evaluation method to delegate certain types of clients to certain CSRs. Together, it was viewed as an important means for front-office workers to self-augment and preserve learning from coworkers for more advanced issues.
@@ 6.3. Pitfalls: Challenges in Integrating Empathetic AI as a Coworker
The paper, so far, has described the utility of an empathetic AI such as Pro-Pilot for on-task ER. Our study design involved a simulation of CSR tasks to probe participants on how AI might provide emotional support through tasks. After the task, the probe served as an anchor for participants to conceive new workflows with such a tool. Their responses indicated the critical limitations in the existing implementation of empathetic AI. We explain these to demarcate the scope of empathetic AI so that we can consider alternative solutions to client incivility.
@@ 6.3.1. On-task empathetic AI has diminishing marginal utility
One of the key aspects of task load is temporal demands (Hart, 2006). We learned that when a CSR is unable to manage client incivility, it leads to a vicious spiral of added burdens (P01, P04, P14, P17, P18, P19): CSR are incentivized to resolve complaints quickly $\rightarrow$ when clients get frustrated, it disrupts their workflow and extends the resolution time $\rightarrow$ the client gets more frustrated as the resolution gets slower $\rightarrow$ the CSR gets further distressed because they start accumulating a backlog of unresolved clients. With this context, participants were wary if actual deployments of empathetic AI like Pro-Pilot might interrupt their efficiency.
@@
P11 pointed out that the external demands of high efficiency interactions can make support insights from Pro-Pilot redundant over time. Similarly, P06 rushed through some of the insights because she found them repetitive. Displaying additional panes could cause information overload. Our structural analysis showed that Pro-Pilot’s messages may need higher reading comprehension (Section 4.2.1). As a solution, P13 recommended that Emo-Reframe should highlight the key takeaway for a quick glance while still maintaining the long-form version for deeper reflection. Based on these experiences, subsequent designs should consider alternatives to the persistent support panes.
@@ 6.3.2. Empathetic AI lacks the social connectedness offered by coworkers
The previous section discussed limitations in the existing social dynamics of coworker support (Section 6.2). Having said that, coworkers remain integral to a CSR’s emotional health. Since they are situated in the same conditions, they are likely to understand the situation better than other social support sources, such as friends (Carmeli et al., 2009). Moreover, participants recognized that building social relationships with coworkers goes beyond work-related support — “You could talk to them anytime, you could see them anytime, talk to them about anything else” (P19). Through our conversations, we distilled that participants felt Pro-Pilot lacks the lived experience of coworkers. Remember, Pro-Pilot’s language rarely used interpersonal pronouns (Section 4.3). Participants perceived their coworkers’ support to be more meaningful because they shared the same experiences (P10, P12, P13, P14, P16). The shared experience convinces the CSR that the advice is more relatable. P13 explained that his coworkers’ advice is valuable because, “I’ve been in a situation like this before, and here’s how I handled it.” P12 even called his conversations with coworkers on the same client as therapeutic. Beyond the shared experience, each coworker also brings in their own unique diverse perspectives (P07). Taken together, it provides the CSR the psychological safety needed to express their emotional concerns and be receptive to support (Carmeli et al., 2009). Enhancing Pro-Pilot with lived experience is non-trivial, but participants could foresee some possibilities.
@@
Above, P12 alluded to the importance of long-term memory (LTM) in improving Pro-Pilot. Prior work shows the importance of (LTM) in AI for mental health (Jo et al., 2024). Introducing LTM to Pro-Pilot could help it learn how a CSR deals with clients and provide more relatable suggestions based on past encounters (P07, P13, P16). Expanding Pro-Pilot’s LTM with experiences of other CSRs could enable AI to mimic coworkers who “flag” emotionally challenging clients and prepare the CSR for encounters (P10). Beyond LTM, another approach to improve Pro-Pilot would be to build a unique model of each CSR’s emotional predisposition. P01 suggested enhancing Pro-Pilot with sensing whereas P13 suggested an option for users to explicitly disclose their objectives to Pro-Pilot. Arguably, human coworkers will be irreplaceable, but some of the contemporary advancements hold promise in replicating experience in AI.
@@ 7. Discussion
Our study with CSRs exhibits the value of emotional support from AI in responding to client incivility.Pro-Pilot presents one of the first applications of LLMs to mitigate intense emotional labor through empathetic human-AI interactions. Emotional labor is not unique to front-office work and our findings are relevant to all sorts of workers, such as information workers (Kuhlthau, 1999), who regularly interact with humans as a part of their job function. The messages from Pro-Pilot distinguish itself from humans and other LLMs. The linguistic analysis already revealed high expressions of empathy and adaptability (Section 4.2.2), but its verbosity, readability, and analytical style were notably different from humans. Yet, the high empathy ratings from CSRs (Section 4.4) indicate that Pro-Pilot’s verbosity might have been associated with thoroughness, readability with compassion, and analytical style with actionability. Subsequently, our simulation exercises revealed the processes through which Pro-Pilot’s emotional component, Emo-Reframe, could help CSRs reduce the demands of uncivil interactions. Despite an additional block of insight, CSRs found Emo-Reframe as helpful as problem-solving insights. Moreover, these studies revealed important opportunities for technological interventions given the current state of social support at work (Section 6). The following discussion aims to anticipate the socio-technical advancements and considerations needed to make empathetic AI-coworkers, such as Pro-Pilot, available in front-office work.
@@ 7.1. Design & Technological Implications
Opportune Moments for On-Task AI Empathy: Clients come in all forms. Actual clients may not always act rude and their emotional expressions, along with its effect on the CSR, are likely to vary by different degrees. Consequently, the need and impact of empathetic support can vary within each interaction. Isolating opportune moments for Pro-Pilot to deliver messages can help reduce the redundancy between messages (Section 4.2.1, Section 6.3.1). The literature on workplace affect sensing (Kaur et al., 2020), stress sensing (Hernandez et al., 2011), and emotional labor sensing (Park et al., 2024) can contribute to making these AI-coworkers act as just-in-time adaptive interventions. Studies on behavioral health recommend modulating the effort required to follow an intervention based on the time of delivery (Jacobson et al., 2022). Depending on the specific context and the CSR’s ability to reflect on additional insight, Pro-Pilot may be trained to produce messages of differing degrees of complexity (Ganapini et al., 2023). Arguably, some of these approaches involve imposing additional sensing into the workers’ ecosystem and we need significant advancements in these studies to accurately model worker’s emotional state in a social context (Das Swain et al., 2022). Alternatively, contemporary research on LLMs shows the possibility of learning worker preferences through their usage patterns (Ning et al., 2024). The next iteration of AI-coworkers needs to be able to anticipate and adapt to CSR needs.
@@
Situated AI-Empathy and Harnessing Experiences: Coworkers also come in all forms. New remote work paradigms need new approaches to social technologies (Das Swain et al., 2020a). It is important for a CSR to perceive their coworker as trustworthy and respectful (Yang and Lau, 2019), even if it is AI (Bartneck et al., 2009).The current iteration of Pro-Pilot did not include any major anthropomorphic aspects to delineate the value of the support messages. Subsequent designs, however, can incorporate more anthropomorphic aspects (e.g., name, appearance, and tonality) to improve the emotional connectedness between the CSR and the AI-coworker (Hermann, 2022). At the same time, personalizing the conversations to users can make them more dynamic and human-like (Section 4.2.1). One way to do this is by learning from the CSR and other CSRs using long-term memory (LTM). Jo et al. found that AI agents that leverage LTM were perceived to be more personal and emotionally supportive (Jo et al., 2024). Beyond personalizing, LTM could also play a role in referring the CSR to real human stories retrieved from its memory. After all, many workers learn “on-the-job,” which is a form of social learning, or learning by observing those in one’s social group (Shteynberg and Apfelbaum, 2013). Future iterations could also be trained on human experiences that workers disclose on public social media (Saha et al., 2019; Das Swain et al., 2020b).. These practices are already common and unique to human coworker social networks (Section 6.3). AI-coworkers could act as a medium to share these experiences between CSRs and help them learn ER through each other.
@@
Scaffolding Emotions in Task-Adjacent Moments: The lack of interventions for on-task cognitive reframing (Slovak et al., 2023) motivated us to design Pro-Pilot as an assistant that can intercept conversations with empathetic suggestions (Section 3). Instead, front-office work incorporates off-task ER methods. Some of these are infrastructural (e.g., training modules) and others are socio-cultural (e.g., internal forum for venting). We believe Pro-Pilot can not only co-exist with these off-task approaches, but it can also help bridge both approaches. For CSRs who regularly interact with the same clients, Pro-Pilot could learn from conversations to create a form of emotional briefing. This early information could act similarly to the usage of AI for planning, collaboration, and communication (Morrison et al., 2024). In turn, it can also be combined with AI for scheduling tasks and ensure workers can focus on important client-interactions when they have maximum resources available (Das Swain et al., 2023b) and also have opportunities to synchronize with their coworkers (Das Swain et al., 2019a). After the task, Pro-Pilot could also provide an emotional debriefing after an intense conversation to help the CSR recover in a way that resembles a coach. Software engineers are already appropriating AI for post-task recommendations (Cabrero-Daniel et al., 2024). What Pro-Pilot learns within the task can be extrapolated in between tasks to help CSRs smoothly ramp in and out of emotional labor.

### v2

@@ 6.1. Utility: Functions of Embedding Empathetic AI into Uncivil Conversations
We found that Emo-Label mirrors those coworkers who simply acknowledge the emotional downsides of the job, without actually providing a way forward. P20 called out the desensitization of his coworkers, “I wouldn’t say they’d be helpful, they see it like it’s a part of my job.” By contrast, participants could identify that Emo-Reframe uniquely varied from simple sentiment classification.
@@ 6.1.1. Empathetic messages help CSRs avoid negative thinking traps
In the face of client incivility, a CSR is likely to feel negatively about the situation. Participants reported learning to dissociate from emotions while on task, but the overwhelming volume of interactions can make complete dissociation challenging (P01, P04, P05, P06, P08, P14, P18, P19, P20). However, as part of their role, a CSR needs to suppress their emotional impulse. P13 described her internal method, “whatever your reaction is, you have to keep that in your head and then decide how you’re going to respond professionally.” While Surface Acting has its benefits (Grandey, 2000), only changing responses at a superficial level contributes to dissatisfaction at work (Côté and Morgan, 2002). Instead, in line with recommendations from emotion literature (Grandey, 2000), we found that Emo-Reframe enabled Deep Acting — adjusting how participants cognitively evaluated their experience.
@@ 6.1.2. Empathetic messages give CSRs an opportunity to recenter their problem solving
Other participants echoed P07’s experience. Essentially, Emo-Reframe helped the participants center their thoughts and continue working towards resolutions. P01 particularly drew a parallel between Emo-Reframe and a coach who would help her maintain “momentum.” Much like how Info-Guide had an affective purpose, participants perceived Emo-Reframe to have a procedural purpose that helped them complete their task. Care-Pilot’s message drew their attention to their objectives and motivated them to work towards it. Note, we designed the empathetic messages with examples of empathy that were high in specificity (Sharma et al., 2023). Participants’ descriptions indicated regaining confidence and psychological safety due to Emo-Reframe. P05 stated, “it feels validating that you’re not necessarily making a mistake.” As a result, participants envisioned their job proficiency improving with empathetic support embedded in the messages.
@@ 6.1.3. Empathetic messages humanize the client
By reading the Emo-Reframe panel, participants were no longer speculating the clients’ perspective. P03 actually became more observant of the client and modified his language to be more accommodating. Participants were more willing to service the clients as “someone alive” (P17). The assistance of Emo-Reframe helped the participants rejuvenate their interest in the client’s emotional state. Not only was this meaningful to their personal wellbeing, but also enables them to build longer-term relationships with clients.
@@ 6.2. Complementarity: Situations where Empathetic AI can Stand-In for Coworkers
Organizations employ multiple CSRs to handle a large volume of client interactions. Regardless of the peculiarity in each client, CSRs together form a “united front” (P04). Coworkers play a key role in supporting CSRs and their empathy is fundamental to mitigating the pressures of client incivility (Ho and Gupta, 2014; Sakurai and Jex, 2012). After interacting with Care-Pilot, our participants compared and contrasted its function with that of their coworkers’. Based on these reflections, we identified two key gaps in existing coworking paradigms of front-office work that emotional support from Care-Pilot can address.
@@ 6.2.1. On-task emotional AI can reduce burden in coworker support
Interactions with clients are nuanced. When the conversation reaches a deadlock, a CSR might want to reach out to their coworkers. However, in cases of incivility, our participants brought to light a double-sided burden problem. Incivility leads to exhaustion (Brotheridge and Grandey, 2002) and can hinder a CSRs ability to seek support. Consider P01’s example, “It’s just hard to bring them into my conversations when I’ll have to explain to them like what is going on.” Support requires some disclosure, but disclosing the complexities of a situation can be burdensome (Bazarova, 2015). Meanwhile, it can also be burdensome for the coworker to be emotionally available for the problem (Fadden et al., 1987). Since the CSR is likely to reach out to a coworker who shares the client, that coworker might be more focused on the task than emotions. By contrast, they viewed Care-Pilot as a tool that can be present throughout their task, observe emotional duress, and provide support immediately. P06 captures the unique opportunity of Care-Pilot’s emotional support:
The Care-Pilot was kind of there with me in the moment, whereas I can’t really have those conversations while I’m trying to have a conversation with the customer. I’ve never had a little emotional support buddy, like that before.” — P06
A CSR might put a client on hold to consult a coworker about procedure, but emotional aspects were discussed after the interaction. Instead, “Care-Pilot is an immediate solution” (P10) for managing emotions of the task. Our lexical evaluation (Section 4.2.2) also confirmed that Care-Pilot was capable of adapting to the specifics of the situation better than human coworkers. Another value proposition of Care-Pilot was that it reduces the need for emotional oversight. P08, who now supervises other CSRs as well, imagined that Care-Pilot would take away the need to “always be on edge, or a constant vigilance, to make sure that my agents … were (not) being insensitive to the client’s question.” With Care-Pilot CSRs stand to reprioritize which emotional discussions they should have with their coworkers.
@@ 6.2.2. Simulating incivility with emotional support can train CSR to use ER
Beyond on-boarding manuals, many CSRs learn on-the-job through other more experienced coworkers (Shteynberg and Apfelbaum, 2013). This learning is hampered because of conflicts that arise among coworkers. If one CSR fails at a task, it is typically escalated to their coworkers. Our participants pointed out that the increased workload of the coworker makes them less inclined to teach the CSR better approaches. This arrangement can strain the social ties between front-office workers. “At work, there are no real friends,” said P04. P01 found herself unable to consult coworkers for her emotional needs because they were not close. P07 even experienced being put down by her coworkers in front of clients. These tensions reduce the ability of early career CSRs to learn from exemplar coworkers. Our study recruited CSRs with at least 2 years of experience. Several expressed their desire to have had a tool like Care-Pilot for training them early.
When they face stuff like this they tend to break down. They need people around them to tell them stuff like what Care-Pilot was doing for me.” — P20
In line with P20’s quote above, P13 — another participant whose role includes supervision — stated that exercising client interactions with Emo-Reframe could help CSRs to develop the ”mindset” needed for client interactions. “ Some of those live reminders, could be really beneficial for folks, who are less practiced or just have more trouble with that kind of emotional regulation,” said P13. Training up the CSRs was seen as an important way to make participants self-reliant. Recent literature shows that workers can outperform their personality based stereotype if they follow healthier behavior patterns (Das Swain et al., 2019b). P18 wants to handle client incivility on her own. With Care-Pilot she felt that her “escalation techniques” would no longer be scrutinized. Some participants saw themselves adapting Care-Pilot into an evaluation method to delegate certain types of clients to certain CSRs. Together, it was viewed as an important means for front-office workers to self-augment and preserve learning from coworkers for more advanced issues.
@@ 6.3. Pitfalls: Challenges in Integrating Empathetic AI as a Coworker
The paper, so far, has described the utility of an empathetic AI such as Care-Pilot for on-task ER. Our study design involved a simulation of CSR tasks to probe participants on how AI might provide emotional support through tasks. After the task, the probe served as an anchor for participants to conceive new workflows with such a tool. Their responses indicated the critical limitations in the existing implementation of empathetic AI. We explain these to demarcate the scope of empathetic AI so that we can consider alternative solutions to client incivility.
@@ 6.3.1. On-task empathetic AI has diminishing marginal utility
One of the key aspects of task load is temporal demands (Hart, 2006). We learned that when a CSR is unable to manage client incivility, it leads to a vicious spiral of added burdens (P01, P04, P14, P17, P18, P19): CSR are incentivized to resolve complaints quickly $\rightarrow$ when clients get frustrated, it disrupts their workflow and extends the resolution time $\rightarrow$ the client gets more frustrated as the resolution gets slower $\rightarrow$ the CSR gets further distressed because they start accumulating a backlog of unresolved clients. With this context, participants were wary if actual deployments of empathetic AI like Care-Pilot might interrupt their efficiency.
@@
P11 pointed out that the external demands of high efficiency interactions can make support insights from Care-Pilot redundant over time. Similarly, P06 rushed through some of the insights because she found them repetitive. Displaying additional panes could cause information overload. Our structural analysis showed that Care-Pilot’s messages may need higher reading comprehension (Section 4.2.1). As a solution, P13 recommended that Emo-Reframe should highlight the key takeaway for a quick glance while still maintaining the long-form version for deeper reflection. Based on these experiences, subsequent designs should consider alternatives to the persistent support panes.
@@ 6.3.2. Empathetic AI lacks the social connectedness offered by coworkers
The previous section discussed limitations in the existing social dynamics of coworker support (Section 6.2). Having said that, coworkers remain integral to a CSR’s emotional health. Since they are situated in the same conditions, they are likely to understand the situation better than other social support sources, such as friends (Carmeli et al., 2009). Moreover, participants recognized that building social relationships with coworkers goes beyond work-related support — “You could talk to them anytime, you could see them anytime, talk to them about anything else” (P19). Through our conversations, we distilled that participants felt Care-Pilot lacks the lived experience of coworkers. Remember, Care-Pilot’s language rarely used interpersonal pronouns (Section 4.3). Recent findings comparing LLM’s social support to that of peers in online communities reflected a similar gap (Saha et al., 2025). Our participants perceived their coworkers’ support to be more meaningful because they shared the same experiences (P10, P12, P13, P14, P16). The shared experience convinces the CSR that the advice is more relatable. P13 explained that his coworkers’ advice is valuable because, “I’ve been in a situation like this before, and here’s how I handled it.” P12 even called his conversations with coworkers on the same client as therapeutic. Beyond the shared experience, each coworker also brings in their own unique diverse perspectives (P07). Taken together, it provides the CSR the psychological safety needed to express their emotional concerns and be receptive to support (Carmeli et al., 2009). Enhancing Care-Pilot with lived experience is non-trivial, but participants could foresee some possibilities.
@@
Above, P12 alluded to the importance of long-term memory (LTM) in improving Care-Pilot. Prior work shows the importance of (LTM) in AI for mental health (Jo et al., 2024). Introducing LTM to Care-Pilot could help it learn how a CSR deals with clients and provide more relatable suggestions based on past encounters (P07, P13, P16). Expanding Care-Pilot’s LTM with experiences of other CSRs could enable AI to mimic coworkers who “flag” emotionally challenging clients and prepare the CSR for encounters (P10). Beyond LTM, another approach to improve Care-Pilot would be to build a unique model of each CSR’s emotional predisposition by leveraging the potential of LLMs to replicate mental health traits (Choube et al., 2024). P01 suggested enhancing Care-Pilot with sensing, whereas P13 suggested an option for users to explicitly disclose their objectives to Care-Pilot. Arguably, human coworkers will be irreplaceable, but some of the contemporary advancements hold promise in replicating experience in AI.
@@ 7. Discussion
Our study with CSRs exhibits the value of emotional support from AI in responding to client incivility. Care-Pilot presents one of the first applications of LLMs to mitigate intense emotional labor through empathetic human-AI interactions. Emotional labor is not unique to front-office work and our findings are relevant to all sorts of workers, such as information workers (Kuhlthau, 1999), who regularly interact with humans as a part of their job function. The messages from Care-Pilot distinguish itself from humans and other LLMs. The linguistic analysis already revealed high expressions of empathy and adaptability (Section 4.2.2), but its verbosity, readability, and analytical style were notably different from humans. Yet, the high empathy ratings from CSRs (Section 4.4) indicate that Care-Pilot’s verbosity might have been associated with thoroughness, readability with compassion, and analytical style with actionability. Subsequently, our simulation exercises revealed the processes through which Care-Pilot’s emotional component, Emo-Reframe, could help CSRs reduce the demands of uncivil interactions. Despite an additional block of insight, CSRs found Emo-Reframe as helpful as problem-solving insights. Moreover, these studies revealed important opportunities for technological interventions given the current state of social support at work (Section 6). The following discussion aims to anticipate the sociotechnical advancements and considerations needed to make empathetic AI-coworkers, such as Care-Pilot, available in front-office work.
@@ 7.1. Design & Technological Implications
Opportune Moments for On-Task AI Empathy: Clients come in all forms. They may not always act rude and their emotional expressions, along with its effect on the CSR, are likely to vary by different degrees. Consequently, the need and impact of empathetic support can vary within each interaction. Isolating opportune moments for Care-Pilot to deliver messages can help reduce the redundancy between messages (Section 4.2.1, Section 6.3.1). The literature on workplace affect sensing (Kaur et al., 2020) and stress sensing (Hernandez et al., 2011; Mattingly et al., 2019) can contribute to making these AI-coworkers act as just-in-time adaptive interventions. Meanwhile, LLMs are exhibiting increasing accuracy in determining mental health labels from text (Xu et al., 2024). Note, however, that designing agent-based wellbeing interventions to user receptivity requires dynamic modeling of users’ motivations (Mishra et al., 2021, 2024). Studies on behavioral health recommend modulating the effort required to follow an intervention based on the time of delivery (Jacobson et al., 2022). Depending on the specific context, mental state, and the CSR’s ability to reflect on additional insight, Care-Pilot may be trained to produce messages of differing degrees of complexity (Ganapini et al., 2023). Commodity devices like smartwatches are becoming increasingly sophisticated for inferring momentary stress (Toshnazarov et al., 2024). Arguably, some of these approaches involve imposing additional sensing into the workers’ ecosystem, and therefore, we need significant advancements in these studies to reliably model worker’s emotional state in a social context (Das Swain et al., 2022). Alternatively, contemporary research on LLMs shows the possibility of learning worker preferences through their usage patterns (Ning et al., 2024). The next iteration of AI-coworkers needs to be able to anticipate and adapt to CSR needs.
@@
Situated AI-Empathy and Harnessing Experiences: Coworkers also come in all forms. New remote work paradigms need new approaches to social technologies (Das Swain et al., 2020a). It is important for a CSR to perceive their coworker as trustworthy and respectful (Yang and Lau, 2019), even if it is AI (Bartneck et al., 2009).The current iteration of Care-Pilot did not include any major anthropomorphic aspects to delineate the value of the support messages. Subsequent designs, however, can incorporate more anthropomorphic aspects (e.g., name, appearance, and tonality) to improve the emotional connectedness between the CSR and the AI-coworker (Hermann, 2022). At the same time, personalizing the conversations to users can make them more dynamic and human-like (Section 4.2.1). One way to do this is by learning from the CSR and other CSRs using long-term memory (LTM). Jo et al. found that AI agents that leverage LTM were perceived to be more personal and emotionally supportive (Jo et al., 2024). Beyond personalizing, LTM could also play a role in referring the CSR to real human stories retrieved from its memory. After all, many workers learn “on-the-job,” which is a form of social learning, or learning by observing those in one’s social group (Shteynberg and Apfelbaum, 2013). These practices are already common and unique to human coworker social networks (Section 6.3). AI-coworkers could act as a medium to share these experiences between CSRs and help them learn ER through each other. Future iterations could also be trained on human experiences that workers disclose on public social media (Saha et al., 2019; Das Swain et al., 2020b). However, in light of suggesting anthropomorphizing, we also caution against it. Research shows that Human-AI interactions with human-like agents can lead to parasocial relationships, where users only have illusory connections to the agent as it is not a real person (Maeda and Quan-Haase, 2024). Such relationships have been shown to have concerning effects on people’s mental health. For instance, users’ overreliance on human-like AI for emotional support can backfire when the agent behaves contrary to their expectations (Laestadius et al., 2024). Given the sensitivity of emotional labor we advise careful examination of emerging social dynamics within HAI as researchers design new AI-assistants for empathetic coworking.
@@
Scaffolding Emotions in Task-Adjacent Moments: The lack of interventions for on-task cognitive reframing (Slovak et al., 2023) motivated us to design Care-Pilot as an assistant that can intercept conversations with empathetic suggestions (Section 3). Instead, front-office work incorporates off-task ER methods. Some of these are infrastructural (e.g., training modules) and others are sociocultural (e.g., internal forum for venting). We believe Care-Pilot can not only co-exist with these off-task approaches, but it can also help bridge both approaches. For CSRs who regularly interact with the same clients, Care-Pilot could learn from conversations to create a form of emotional briefing. This early information could act similarly to the usage of AI for planning, collaboration, and communication (Morrison et al., 2024). In turn, it can also be combined with AI for scheduling tasks and ensure workers can focus on important client-interactions when they have maximum resources available (Das Swain et al., 2023b; Hernandez et al., 2024) and also have opportunities to synchronize with their coworkers (Das Swain et al., 2019a). After the task, Care-Pilot could also provide an emotional debriefing after an intense conversation to help the CSR recover in a way that resembles a coach. A practical way forward would be to augment Care-Pilot with LLM tools that encourage self-reflection of stressful experiences (Song et al., 2025). Software engineers are already appropriating AI for post-task recommendations (Cabrero-Daniel et al., 2024). What Care-Pilot learns within the task can be extrapolated in between tasks to help CSRs smoothly ramp in and out of emotional labor.